OpenStudy (anonymous):

Twenty times the square of a non-zero number is equal to 20 times the number ! Whta is the number ? :) lets see if you guys can do it ? :D

OpenStudy (bahrom7893):

20x^2 = 20x 20x = 20 x = 1

OpenStudy (anonymous):

x=0 :|

OpenStudy (anonymous):

you divided out a soln

OpenStudy (bahrom7893):

Non-zero number...

OpenStudy (anonymous):

good :) you are correct :P and thanks for putting the equation :) how about this ten times the square of a non zero number is equal to seventy time the number ? and how to solve >?

OpenStudy (bahrom7893):

0 is obvioulsy the solution.. 20x^2-20x = 0 20x(x-1) = 0, and one is 0 and the other one is 1

OpenStudy (anonymous):

do you use to same equation ?

OpenStudy (bahrom7893):

no that's not for that... wait

OpenStudy (anonymous):

0 was wronge ; so 1 was the correct option ; so what is x ? the number your looking for am i correct ?

OpenStudy (bahrom7893):

ten times the square of a non zero number is equal to seventy time the number ? 10*x^2 = 70x

OpenStudy (anonymous):

oh :( sorry im super gullible .

OpenStudy (bahrom7893):

10x^2-70x=0 10x(x-7)=0 x=0, but u dont want that since it says non zero number x-7 = -, so x=7

OpenStudy (anonymous):

thanks love :) i kind of get what im doing now . thanks soo much :)

OpenStudy (bahrom7893):

np

OpenStudy (anonymous):

how old are you that your so smart :o im 17 if you wanna know :)

OpenStudy (bahrom7893):

im 17

OpenStudy (anonymous):

really :D ? thats cool . are you lieing ? lol

OpenStudy (anonymous):

*lying .

OpenStudy (bahrom7893):

if u dont wanna believe me its fine, i dont care..y would i lie about my age..

OpenStudy (anonymous):

lol alright im just kidding :P thanks again though .

OpenStudy (anonymous):

do you have a facebook ? lol :) you should add me :) lol

OpenStudy (bahrom7893):

i think im through adding people from math websites on fb..my wall was full of random problems until i deleted them..

OpenStudy (anonymous):

lmaoo; that sucks ; well as you can see imn not in to math . :D but i totally get that lmaooo i would of done the same thing :P well im from jersey :) and not that smart lol .

OpenStudy (anonymous):

* well im not into it but i love learning it idk why lol .